HUBBLE SPACE TELESCOPE DAILY REPORT #1196 PERIOD COVERED: 0000Z Daily Status Report as of 217/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 1.2 Completed Three Sets of WF/PC-2 5091 (GTO High Latitude Parallel Program: Cycle 4) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations greater than +/-30 degrees. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 High, High Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were completed as planned, and no problems were eported. 1.4 Completed FOS/RD 5529 (Cycle 4 FOS Discriminator Test) The Faint Object Spectrograph (Red Side) was used to run discriminator tests to determine if the discriminator settings need adjustment from the Cycle 3 Values. Dark, Internal Flat, and Wave calibrations observations were taken. The Observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOC/96 5522 (FOC Post-COSTAR Polarization Calibration) The Faint Object Camera (f/96) was used to make polarization calibration observations using standard object NGC-5272 as a source to help fine the transmission coefficients. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 5569 (WF/PC-2 Cycle 4 Calibration: Internal Flats) The WF/PC-2 was used to make flat field calibration images with the VIS lamp, and the internal source. The observations were completed as planned, but a small segment was lost due to a taperecorder track change. 1.7 Completed FOC/96 5526 (FOC UV Throughput Monitoring: Cycle 4) The Faint Object Camera (f/96) was used to make calibration observations using BPM16274. HST underwent 1 guide star recentering(s) for a total of 5 seconds. During the recentering the interval HST's pointing was controlled by gyros, which allow a small amount of drift during the interval. 1.8 Completed Two Sets of GHRS/1 5595 (GHRS ECH-A Wavelength Monitor: Cycle 4) The High Resolution Spectrograph was used to make wavelength spectral lamp calibration observations to measure the instrumental characteristics of the Echelle-A spectrograph. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Two Sets of WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of NGC-2985 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 5563 (WF/PC-2 Cycle 4 Photometric Calibration Monitor 1: UV /OPT STD.) The WF/PC-2 was used to make photometric calibration images of ultraviolet standard star GRW+70D582 and monitor the photometric stability and quantum efficiency of the WF/PC-2. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled :13 Successful :13 2.3 Significant HSTARs: HSTAR 4863: documents a suspected problem with some element of the ground software. The events printout shows a termination of the telemetry link to the ESS at the same time data was acquired after a switch to 32k telemetry. 2.4 Operations Notes: A small adjustment in the COSTAR's FOC M1 mirror in X tilt was made on 215/1400z to remove residual coma in the FOC PSF. The result will be more symmetrical PSF, especially in the UV. The move was a single motor step in the negative direction (1 motor step = 7.2 a-s in tilt). Due to hysteresis within the mechanism the commanding required to achieve a single negative motor step was -46 steps, +45 steps. On day 215 at approx. 04:28 the WFPC-2 science data interface hung due to a PMDB editting error which caused infrequently used WFPC shutter commands to overlap into a WFPC science data read-out. These infrequently used commands slowed the read-out of the WFPC CCDs. Since WFPC command execution is synched with the major frame a commanding an entire minute. The delay in the following read-out, causing the WFPC interface to hang. This ultimately caused a SDF lockout. The SDF and WFPC science data interface were recovered via r/t commanding on 215/1800z. ********************** ********************** 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /Dave L. -END-